OpenStudy (anonymous):

Find the area bounded by the curve x=t-1/t, y=t=1/t, y=65/8

OpenStudy (astrophysics):

Is that \[x= t - \frac{ 1 }{ t }~~~y=t-\frac{ 1 }{ t }~~~y=\frac{ 65 }{ 8 }\]

OpenStudy (ali2x2):

@Astrophysics Yes it looks like it

OpenStudy (ali2x2):

y=t=

hartnn (hartnn):

y could be t+1/t since, + and = are on the same key of the keyboard

OpenStudy (astrophysics):

I think you're right hartnn! So areas in parametric curves we have \[\int\limits_{a}^{b}g(t)f'(t) dt\]

OpenStudy (astrophysics):

 now to see where the curve touches the line we set \[t+\frac{ 1 }{ t } = \frac{ 65 }{ 8 }\] solving for t will tell us at which time our curve will touch the line, then you can make a graph to make your curve. So we get our time to be \[t=\frac{ 1 }{ 8 } ~~~and ~~~ t = 8\]

OpenStudy (astrophysics):

To find our values for x, you can make a graph as I suggested (just too keep your values and see what's going on), or you can just do when t = 1/8 \[x= t-\frac{ 1 }{ t } \implies \frac{ 1 }{ 8 } - \frac{ 1 }{ \frac{ 1 }{ 8 } } = ...\] \[x= 8-\frac{ 1 }{ 8 } = ...\] You'll then have your x,y,t values and you should be able to graph it, that's the major part, then setting up the integral shouldn't be too bad.

OpenStudy (astrophysics):

I don't plan on doing this whole thing for you, so if you're there, please ask for help.

ganeshie8 (ganeshie8):

Alternatively, if you don't like parametric curves, you may try eliminating the parameter easily by noticing that \[(a+b)^2-(a-b)^2 = 4ab\]

ganeshie8 (ganeshie8):

\[y^2 - x^2 = \left(t+\frac{1}{t}\right)^2 - \left(t-\frac{1}{t}\right)^2=4*t*\frac{1}{t} = 4 \]

ganeshie8 (ganeshie8):

so the problem is equivalent to finding the area between curves : \(y^2-x^2=4\) and \(y=65/8\)

OpenStudy (astrophysics):

So much easier xD

ganeshie8 (ganeshie8):

Not really, looks both require almost same effort haha!

OpenStudy (astrophysics):

Ah, I see, I just think it's easier to see the curve with your method...but yeah I think it's equal effort hehe.

OpenStudy (anonymous):

still not sure how to get the right answer.

OpenStudy (astrophysics):

Ok well, doing it the method I suggested, what values did you get for x?

OpenStudy (astrophysics):

\[x= t-\frac{ 1 }{ t } \implies \frac{ 1 }{ 8 } - \frac{ 1 }{ \frac{ 1 }{ 8 } } = ...\] and \[x= 8-\frac{ 1 }{ 8 } = ...\]

OpenStudy (astrophysics):

We'll go over this step by step, so just find what x's are and we can continue :)